in a triangle ABC, angle b is 5 times angle A and angle C is 4 degrees less than 4 times angle A. what is the size of the angles?
Answer by dabanfield(803)

in a triangle ABC, angle b is 5 times angle A and angle C is 4 degrees less than 4 times angle A. what is the size of the angles?

We know by a theorem that the sum of the angles of a triangle equals 180. Let x be angle A. Then we have:
Angle A = x
Angle B = 5*x
Angle C = 4*x - 4
The sum of these must be 180 so:
x + 5*x + (4*x - 4) = 180
Solve the above for x then calculate 5*x and 4*x - 1.